OpenStudy (anonymous):

An equation is shown below: 6(2x – 11) + 15 = 3x + 12 Part A: Write the steps you will use to solve the equation, and explain each step. (6 points) Part B: What value of x makes the equation true? (4 points)

OpenStudy (kidrah69):

Ok so do you know how to solve equations?

OpenStudy (anonymous):

kind of

OpenStudy (kidrah69):

Ok show me what you know :)

OpenStudy (anonymous):

Im just really stuck on this problem and i dont even know where to begin

OpenStudy (kidrah69):

Think of Order of operations :)

OpenStudy (kidrah69):

Scratch that >_< first distribute.

OpenStudy (anonymous):

Well im guessing we have to start by solving whats inside the paraenthensis

OpenStudy (anonymous):

oh never mid

OpenStudy (kidrah69):

You figured it out?

OpenStudy (anonymous):

no i still need help :)

OpenStudy (kidrah69):

oh ok :) 1) Distribute

OpenStudy (anonymous):

so im guessing we had 6 and 15 together?

OpenStudy (anonymous):

add

OpenStudy (kidrah69):

no no no we want to MULTIPLY!! :D \[6(2x – 11) + 15 = 3x + 12 \] Think: \[6\times2x=?\] \[6\times-11=?\]

OpenStudy (anonymous):

6 x -11 = -66 and the other one equals 12x

OpenStudy (kidrah69):

yes! :D now we have.... \[12x-66+15=3x+12\] Combine like terms :) -66+15=?

OpenStudy (anonymous):

-51

OpenStudy (kidrah69):

good \[12x−51=3x+12\] now combine like terms! (this is the fun part!) 12-3=? 12+51=?

OpenStudy (anonymous):

ok so it woud euqual 9x= -39

OpenStudy (kidrah69):

nooo.... 12+51=?

OpenStudy (anonymous):

63

OpenStudy (kidrah69):

good what is 63/9?

OpenStudy (anonymous):

7

OpenStudy (kidrah69):

1 )Expand 12x−66+15=3x+12 2 )Simplify 12x−66+15 to 12x−51 12x−51=3x+12 3 )Add 51 to both sides 12x=3x+12+51 4 )Simplify 3x+12+51 to 3x+63 12x=3x+63 5) Subtract 3x from both sides 12x−3x=63 6) Simplify 12x−3x to 9x 9x=63 7 )Divide both sides by 9 x=639 8) Simplify 63/9 to 7 x=7

OpenStudy (kidrah69):

That answers part A and part B also good job! :D

OpenStudy (anonymous):

Thank you so much!

OpenStudy (kidrah69):

No problem! Also welcome to openstudy!! :D